22 June 2013 LBC IQ Technical Observing Night 1

OSA: Steve Allanson
Observatory Manager: John Little
Observers: Steven Villanueva, Olga Kuhn

Summary

It took a while to get started due to the LBC problems and high winds. After opening, we obtained pupil images with MODS, with LBC, all with a 1.5 foot long cardboard extended over the primary. It was not sufficiently long to see obscuration.

We then removed it and checked the LBC Blue/Red rotator centers. These did not change much, ~5 pixels from the header values, ~1".

We did not stay with the plan to collect data for a collimation model as it looked like the seeing was quite good and it was already late - the time needed to make 4 sweeps in elevation 5-6 hours.

We instead obtained several superfoc sequences. We wanted to obtain tech chip data, but not to guide or to send focus corrections, and I tried to adjust the configuration files to accomplish this:
  • Trackers.MinimumScientificExposureTime changed from 64 -> 5 (this change had been made yesterday in anticipation).
  • Trackers.LoopStartDelay changed from 10 -> 1 second (this change never showed up in the lbc.log so I don't know why)
  • Trackers.AzTrackingGain and Trackers.ElTrackingGain both changed from 1.0 -> 0.0 (according to the log, this took effect from 08:00-10:00)
  • Trackers.MinGuidingCorrection was even changed at one point from 0.01 to 1000 but according to log, this change never took effect.
  • Trackers.TypicalExposureTime from 8->4 (never took effect according to the log).
After each change, we stopped/restarted LBCs (we could have run lbckill/lbcstart only, I later remembered), so I am not sure why some changes did not take effect. Likely I was tired and not doing these changes as systematically as we thought?? Remains a mystery... Possibly due to this, we did not have any success obtaining a pair of tech chip images for the 10-sec exposures of the superfoc7 sequences, however we guided during the extra-focal pupil images which was not desired. For the 15-sec test exposure, we obtained tech chip images on one side but not on the other side. And, even if tech chip images are obtained, and guiding is not activated, I did not test the turning off/on of the AO (through a TRUE/FALSE flag). While it would be useful to obtain tech chip images throughout the superfoc sequences, I think more work on this will likely wait for a poor weather night. Original configuration files are restored.

On the superfoc sequences done in good seeing, the best-focus on the blue side differed by 60 microns, from +0.03 to -0.03 for a second set. We have to measure best-focus for all of these images still. See some preliminary best-focus measurements at end of log.

The seeing did not remain so stable for filter focus offsets so we obtained only one other pair, at PA=0 and PA=180 this time, before moving to some tests for rotational trailing. We saw rotational trailing for the OB with the large dithers. We did not see it for an OB with a 0,0 first dither. But I don't know if this is correlated or because the field had just transited and was now setting. Tried another field that was rising.

Plan is to spend the next two nights collecting data for a collimation model and more superfoc sequences to measure filter focus offsets and tilt.

Log

noticed on Thursday afternoon and ongoing until evening LBC Blue Profibus errors. Traced to a dangling fiber and damaged fiber connection. John Little and Steve Allanson repaired this. LBC is working now. Winds were gusting too high to open at the start.

05:00 Open and Steve A doing pointing/collimation

05:13 MODS focus/collimated (roughly). Delta Z = -0.2mm b =1 r=1 pupil radius is ~ 17 pixels. And for Delta Z = -0.4mm b=2 r=2. Pupil radius ~ 31 pixels.

05:25 Attached a rectangular cardboard to the outside of M1, away from the C-rings and near the mirror ventilation hoses. The cardboard is about 1.5 ft long and hopefully enough to be seen in the pupil images.

05:30 Now going back to a BS 9169 (gs 2) to focus/collimate. A lot of Z11 and pincushion on the WFS images - mirror not thermalized yet. Seeing looks like it will be <~ 1".
  • Collimated. Now disable reporting of guiding and Zernike corrections & send a global offset of Z = -0.4 to M1. We did not adjust the guider-instrument offset, just are taking an unguided image. Should we try this while guiding? Exposures are relatively short and we don't do this with LBC, so I think no... but pupils (1-sec) look fuzzy.
  • b=3 r=3 dZ=-0.4mm 1-sec
  • b=4 r=4 dZ=-0.8mm 30-sec (now I think we can discern the swing arm and the cardboard, but we are making the pupil still bigger).
  • b=5 r=5 dZ=-1.2mm 30-sec (clearly see swing arm but not sure about cardboard... it should be ~ 60 deg from base of swing arm (i.e. not at right angles to it)
  • b=6 r=6 dZ=-1.6mm 30-sec
  • b=7 r=7 dZ=-1.6mm 60-sec (really hard to discern the cardboard - we think we know where it is, but not sure...).
  • b=8 r=8 dZ=+1.6mm 60-sec (still hard to discern...)
  • Attaching jpg of mods pupil image. Some fuzzy edge is seen also.

06:00 Switch to LBC. Will take a couple extra/intral focal pupil images, but then will take off the cardboard and obtain data for the collimation model.

06:20 Slew to first LBC target, copoint star. focus/collimate. FPIA estimated seeing 1.4" blue and 0.8" red. At about 2:30 cw from the swing arm there is some obsc on the blue pupil, but not conclusive.

06:27 Run RB_rV100exin4 pupils. Not very convincing --- need longer piece of cardboard.
  • Blue pupils show a "sunburst" around the edges of extra-focal pupils. On MODS as well as LBC. As much as I stretch the contrast on the LBC Red extra-focal pupils, I do not see this effect. _Attaching two jpgs of the set of 4 LBC pupil images (below). One at low contrast, to see the spider, and one at high contrast to show the edge.
Blue extra/intra focal pupils (100 pix diameter):
lbcb.20130622.062545.fits[2] FOCUS -1.320
lbcb.20130622.062653.fits[2] FOCUS 1.280

Red extra/intra focal pupils (100 pix diameter):
lbcr.20130622.062539.fits[2] FOCUS -1.264
lbcr.20130622.062645.fits[2] FOCUS 1.336

06:30 Removing cardboard... it was actually a little closer to 90 deg than we thought.

start taking data for collimation model

06:39 EL 70.
  • dofpia, seeing ~ 1" on dimm,
  • lbcrangebal B=64801 R=64759
MOUNT: delta_IE = -5.52 arcsec, delta_CA = -8.76 arcsec
M1 SX: dX = 0.01 mm, dY = 0.12 mm, dRX = 2.64 arcsec, dRY = -0.26 arcsec 
M1 DX: dX = -0.03 mm, dY = -0.06 mm, dRX = -1.32 arcsec, dRY = 0.73 arcsec 
  • run copoint OB again B 065357 R 065354

rotcenter check
  • B 65652 R 65649 pa=0 and B 70001 R 65958 (images are not close enough to -179.568 deg apart)
  • B 70525 R 70522
Blue (old: 1035, 2924)
 
      Xc=1030.82 Yc=2928.1 from geomap.
      Xc=1031.33 Yc=2928.07 from averages.
Red (old: 1078m 2913)
      Xc=1080.49 Yc=2910.42 from geomap.
      Xc=1080.14 Yc=2910.57 from averages.

Within 5 pixels.

How low elev can we copoint? (not answered --- we saw drift and noticed that the conf file changes made earlier caused us to guide)
  • Elev=18:46
  • We see a lot of drift and notice we are still guiding... Had changed TrackersMinimumScientificExposureTime to 5-sec (from 40) so we can obtain tech chip images during the exposure. Could change also MinGuidingCorrection. from 0.01 to 1000. Still lbc.log file says we are guiding, as well as LBTplot.
  • Changed Trackers.Az/ElTrackingGain to zero. Turned off/on LBCs.
  • FPIA built up a lot of spherical.
  • Examined series of pupil images during next afternoon. Initial pupils do have a dark inner annulus, apparently interpreted by FPIA as spherical. There is drift. We were guiding on left side, but not on right. But max guiding offset was only ~1 pixel (0.26") (grep Xarcs /lbt/log/2013/06/20130622.log or see guidelbc.sx.20130622lbto.tmp). This does not explain the jumps from image to image. Could it be pointing offsets, but these (according to grep AdjustPointing 20130622.events) are subarcsec?. Not sure if PCS applies them... I see messages like
=pcs.command.applyRBTipTilt.warning right PCS ApplyRBTipTilt Warning: application of PSF range balance tip/tilt request is disabled on side right or in MONOCULAR mode= 
throughout the night. I see same messages on INAF science night 20130612, however. Could be coma (z7,z8). How does delta_z7(nm) -> delta_az/el(")? 21.454" dRY , corresponds to -1mm dX, and at FP?

So I don't think guiding or tt corrections were signfiicant contributors to difficult convergence and spurious Z11 and/or Z22 driven in. These may be pupil images to examine for a dependence of FPIA on seeing. Seeing ~ 1.5"

 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  3252     45   -123   -414   -110    600     14    y  1.69  1.35                      /newdata/lbcb.20130622.072153.fits  V-BESSEL
 -1935    101   -184     17      4    471     53    y  1.71  0.82                      /newdata/lbcb.20130622.072329.fits  V-BESSEL
 -1799    114    -86     72     18    440     39    y  1.61  0.77                      /newdata/lbcb.20130622.072502.fits  V-BESSEL
 -1165     52    -76    -10      5    296     14    y  1.57  0.50                      /newdata/lbcb.20130622.072637.fits  V-BESSEL
 -1252     30   -105     92    -62    201     30    y  1.51  0.52                      /newdata/lbcb.20130622.072811.fits  V-BESSEL
  -553     46    -86    -43     26    146     15    y  1.56  0.25                      /newdata/lbcb.20130622.072945.fits  V-BESSEL
 -1043     75    124   -189    -69    192     25    y  1.50  0.45                      /newdata/lbcb.20130622.073119.fits  V-BESSEL
  1095      0      0      0      0   -625      0    y  0.00  0.00                     dofpia_sph3_adjustment             


  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1186    -89     13     23    290    343     23    y  1.33  0.54                      /newdata/lbcr.20130622.072148.fits  r-SLOAN 
 -1377    -89     41    -26     23    328     42    y  1.54  0.59                      /newdata/lbcr.20130622.072326.fits  r-SLOAN 
 -1143   -131    -61    136    -36    269     27    y  1.41  0.49                      /newdata/lbcr.20130622.072459.fits  r-SLOAN 
 -1296    -71     34   -146      3    222      8    y  1.31  0.54                      /newdata/lbcr.20130622.072631.fits  r-SLOAN 
  -584    -64      5    163    -22    196      4    y  1.33  0.28                      /newdata/lbcr.20130622.072808.fits  r-SLOAN 
  -736    -66    -41     97     -1    158      9    y  1.28  0.32                      /newdata/lbcr.20130622.072942.fits  r-SLOAN 
  -612    -60     37   -126     76     90      6    y  1.30  0.27                      /newdata/lbcr.20130622.073116.fits  r-SLOAN 
   929      0      0      0      0   -530      0    y  0.00  0.00                     dofpia_sph3_adjustment             

We don't have the ~ 6hr to collect collim model data, but seeing is good at higher elevations, so tilt/focus offset data instead

08:10 Slew to field at 70 deg elevation. dofpia is taking about z11, z22, but we stop and ask Steve to clear active optics for good measure.

08:15 Active optics cleared, dofpia started again.
  • Finding lots of pupils (~8)
  • There IS a temperature gradient across the mirror. Both blue/red show it.
  • dofpia converged. Run RB_rV_superfoc7.ob (bestfoc 0.027 mm)
lbcb.20130622.082553.fits[2] RB_rV_superfoc7 -0.200 FOCUS
lbcb.20130622.082642.fits[2] RB_rV_superfoc7 -0.140 FOCUS
lbcb.20130622.082732.fits[2] RB_rV_superfoc7 -0.080 FOCUS
lbcb.20130622.082821.fits[2] RB_rV_superfoc7 -0.020 FOCUS
lbcb.20130622.082910.fits[2] RB_rV_superfoc7 0.040 FOCUS
lbcb.20130622.082959.fits[2] RB_rV_superfoc7 0.100 FOCUS
lbcb.20130622.083049.fits[2] RB_rV_superfoc7 0.160 FOCUS
lbcr.20130622.082551.fits[2] RB_rV_superfoc7 -0.144 FOCUS
lbcr.20130622.082631.fits[2] RB_rV_superfoc7 -0.084 FOCUS
lbcr.20130622.082711.fits[2] RB_rV_superfoc7 -0.024 FOCUS
lbcr.20130622.082751.fits[2] RB_rV_superfoc7 0.036 FOCUS
lbcr.20130622.082831.fits[2] RB_rV_superfoc7 0.096 FOCUS
lbcr.20130622.082911.fits[2] RB_rV_superfoc7 0.156 FOCUS
lbcr.20130622.082953.fits[2] RB_rV_superfoc7 0.216 FOCUS
  • No tech chip images obtained. Changed LoopStartDelay from 10 to 1 sec. Turn LBC off/on.

08:50 Slew again to 70 deg field.
  • were pretty well collimated earlier. So directly run lbcrangebal
  • dofpia (converged with seeing 1.14 on blue but 0.81 on red. Still at Temp gradient across both primaries. Ambient temperature decreasing a little.
  • RB_rV_superfoc at PA=0 (bestfoc -0.034 mm)
lbcb.20130622.085823.fits[2] RB_rV_superfoc -0.100 FOCUS
lbcb.20130622.085908.fits[2] RB_rV_superfoc -0.060 FOCUS
lbcb.20130622.085951.fits[2] RB_rV_superfoc -0.020 FOCUS
lbcb.20130622.090035.fits[2] RB_rV_superfoc 0.020 FOCUS
lbcb.20130622.090119.fits[2] RB_rV_superfoc 0.060 FOCUS
lbcr.20130622.085819.fits[2] RB_rV_superfoc -0.044 FOCUS
lbcr.20130622.085855.fits[2] RB_rV_superfoc -0.004 FOCUS
lbcr.20130622.085930.fits[2] RB_rV_superfoc 0.036 FOCUS
lbcr.20130622.090005.fits[2] RB_rV_superfoc 0.076 FOCUS
lbcr.20130622.090039.fits[2] RB_rV_superfoc 0.116 FOCUS

09:14 dofpia again on 70 deg field (now at elev 67 deg)
  • Definitely see a "skirt" on blue pupils --- see 091451
  • Run superfoc7 pa=0 (best-focus -0.02, but there is noise in the plot. DIMM seeing now 1.3" ...)
lbcb.20130622.091705.fits[2] RB_rV_superfoc7 -0.200 FOCUS
lbcb.20130622.091754.fits[2] RB_rV_superfoc7 -0.140 FOCUS
lbcb.20130622.091842.fits[2] RB_rV_superfoc7 -0.080 FOCUS
lbcb.20130622.091930.fits[2] RB_rV_superfoc7 -0.020 FOCUS
lbcb.20130622.092020.fits[2] RB_rV_superfoc7 0.040 FOCUS
lbcb.20130622.092107.fits[2] RB_rV_superfoc7 0.100 FOCUS
lbcb.20130622.092157.fits[2] RB_rV_superfoc7 0.160 FOCUS
lbcr.20130622.091705.fits[2] RB_rV_superfoc7 -0.144 FOCUS
lbcr.20130622.091745.fits[2] RB_rV_superfoc7 -0.084 FOCUS
lbcr.20130622.091824.fits[2] RB_rV_superfoc7 -0.024 FOCUS
lbcr.20130622.091906.fits[2] RB_rV_superfoc7 0.036 FOCUS
lbcr.20130622.091946.fits[2] RB_rV_superfoc7 0.096 FOCUS
lbcr.20130622.092025.fits[2] RB_rV_superfoc7 0.156 FOCUS
lbcr.20130622.092105.fits[2] RB_rV_superfoc7 0.216 FOCUS
    • cannot get tech chip images for these 10-sec exposure times
  • Run superfoc7 pa=180
lbcb.20130622.092417.fits[2] RB_rV_superfoc7_pa -0.200 FOCUS
lbcb.20130622.092506.fits[2] RB_rV_superfoc7_pa -0.140 FOCUS
lbcb.20130622.092555.fits[2] RB_rV_superfoc7_pa -0.080 FOCUS
lbcb.20130622.092645.fits[2] RB_rV_superfoc7_pa -0.020 FOCUS
lbcb.20130622.092734.fits[2] RB_rV_superfoc7_pa 0.040 FOCUS
lbcb.20130622.092823.fits[2] RB_rV_superfoc7_pa 0.100 FOCUS
lbcb.20130622.092912.fits[2] RB_rV_superfoc7_pa 0.160 FOCUS
lbcr.20130622.092415.fits[2] RB_rV_superfoc7_pa -0.144 FOCUS
lbcr.20130622.092455.fits[2] RB_rV_superfoc7_pa -0.084 FOCUS
lbcr.20130622.092534.fits[2] RB_rV_superfoc7_pa -0.024 FOCUS
lbcr.20130622.092616.fits[2] RB_rV_superfoc7_pa 0.036 FOCUS
lbcr.20130622.092657.fits[2] RB_rV_superfoc7_pa 0.096 FOCUS
lbcr.20130622.092737.fits[2] RB_rV_superfoc7_pa 0.156 FOCUS
lbcr.20130622.092817.fits[2] RB_rV_superfoc7_pa 0.216 FOCUS
  • fast OB with 15-sec exp time. Did take a blue tec exp, but not a red one. Tried again and took neither blue nor red tech chip image.

09:43 Restore configuration files. Going to run rotator tracking test.
  • focus/collimate and copoint on field at 80 deg elevation.
  • copoint B 095419 R 095417
  • load and run original dither OB.
    • dithoffx = -700, -350, 0, 350, 700 and dithoffy = -140, -70, 0, 70, 140
  • Guiding corrections are all zeroes!!! But the red and blue channel conf files were restored. Turn off LBCs. lbckill/start and turn back on LBCs.
  • Exposures completely saturated in original exptime 250sec, even at V-BESSEL. Edited OBs to reduce exptime to 150s and change from V/r to S/V.
    • repeat original dither set...(guiding OK now). see no sign of rotational trailing and guider offsets do not look so strange. We are at 77 deg elevation.
  • Go to field that is rising rather than this one that is setting.
    • dofpia
    • lbcrangebal
    • run script which uses original dithers (see above). b 103740 shows some rotational trailing - not obvious but there. r 103728 also shows it.
    • now run script with 0,0 first dither
      • I don't see trailing, but this field is now setting, though still at elev~83. stopped OB after 3/5 dithers
    • going to one more field that is rising and will run 0,0 dither OB:
      • (skip copointing --- we did it already at this elevation a while ago, and it is getting light).
      • run OB which uses 0,0 first dither. dithoffx = 0, -350, 0... and dithoffy=0,-70, 0...

Saturday afternoon

Measured best-focus at center of chip 2 for the 4 blue/red superfoc sequences obtained. I used iraf.obsuit.starfoc with the settings listed below, and I used a 'random' selection of stars near the central ~ 7' x 5' region. I reordered the list to put the middle image (supposedly at best-focus) first, so star selection was based on this close-to-focus image. Dofpia had been run immediately before each of the sequences: a; b and c; but not between sequences c at pa=0 and pa=180.

filename image sequence pa best-focus[mm] FWHM at best-focus [pix] comments
bfoc0a.list 082553-083049 0 0.026 3.87  
bfoc0b.list 085823-090119 0 -0.041 3.76  
bfoc0c.list 091705-092157 0 0 3.79 point at focus 0.04 was above-curve...seeing blew-up?. Deleted all points at this focus from fit. Note this deviation not seen on simul red seqeunce
bfoc180c.list 092417-092912 180 0.025 3.50  
filename image sequence pa best-focus[mm] FWHM at best-focus [pix] comments
rfoc0a.list 082551-082953 0 0.025 4.36 high fwhm, repeated and got best-foc 0.0295 at FWHM 4.21
rfoc0b.list 085819-090039 0 0.026 3.67  
rfoc0c.list 091705-092105 0 0.034 3.25  
rfoc180c.list 092616-092817 180 0.039 3.22  

Repeated blue measurement with smaller region (zoom 1/2 or about 2.8' x 1.8') about center.
filename image sequence pa best-focus[mm] FWHM at best-focus [pix] comments
bfoc0a.list 082553-083049 0 0.028 3.83  
bfoc0b.list 085823-090119 0 -0.032 4.14  
bfoc0c.list 091705-092157 0 0 3.8 still point at focus 0.04 was above-curve... with this focus point, best-focus was -0.02, without it, best-focus ~ 0.005
bfoc180c.list 092417-092912 180 0.028 3.36  

-- OlgaKuhn - 22 Jun 2013
I Attachment Action Size Date Who Comment
blue2013-06-22.LogLog blue2013-06-22.Log manage 8 K 23 Jun 2013 - 21:26 OlgaKuhn blue2013-06-22.Log (FPIA log)
lbcpupils_highcontrast.jpgjpg lbcpupils_highcontrast.jpg manage 48 K 23 Jun 2013 - 02:37 OlgaKuhn LBC extra/intra 100-pix diameter pupil images stretched to illustrate "burst" around blue pupil edge
lbcpupils_lowcontrast.jpgjpg lbcpupils_lowcontrast.jpg manage 34 K 23 Jun 2013 - 02:36 OlgaKuhn LBC extra/intral focal 100-pix diameter pupils
modspupils.jpgjpg modspupils.jpg manage 64 K 23 Jun 2013 - 02:42 OlgaKuhn mods extra (5 and 7) and intra-focal (8) images.
red2013-06-22.LogLog red2013-06-22.Log manage 8 K 23 Jun 2013 - 21:26 OlgaKuhn red2013-06-22.Log
Topic revision: r6 - 24 Jun 2013, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback